# Stationary Policies are Optimal in Risk-averse Total-reward MDPs with EVaR

### v1

@@ 4.2 Algorithms
We now propose a simple algorithm for computing a $\delta$-optimal EVaR policy described in Algorithm 1. The algorithm reduces the EVaR optimal policy to a sequence of solving ERM-TRC problems. As Theorem 4.1 shows, there exists a $\delta$-optimal policy such that it is ERM optimal for some $\beta\in\mathcal{B}(\beta_{0},\delta)$. It is, therefore, sufficient to compute an ERM-optimal policy for one of those $\beta$ values.
Data: MDP and desired confidence $\delta>0$
@@
An advantage of the approach taken by the algorithm is that the overhead of completing optimal solutions for all $\alpha$ values is negligible if one selects an appropriate set $\mathcal{B}$.
@@ 5 Numerical Evaluation
In this section, we illustrate the results of Section 3 and the influence of the risk parameter $\beta$ of the ERM, and discuss the influence of the risk level $\alpha$ of EVaR on tabular transient MDPs that include positive rewards and negative rewards.
First, Figure 2 illustrates the dramatic difference between the discounted criterion and TRC with the ERM criterion for $\beta>0$. We use the discounted and transient MDPs in Figure 1 with parameters $r=-0.2,\,\gamma=0.9,\,\epsilon=0.9$. In the risk-neutral case, they have identical value functions ($\beta=0$); see Appendix B. However, the discounted return is unaffected by $\beta$ because the MDP is deterministic. However, the ERM-TRC return decreases with an increasing $\beta$ and becomes unbounded for $\beta\geq 0.526$. For large $\beta$ values, the return is unbounded because $\rho(\bm{B})\geq 1$ for the corresponding $\bm{B}$ defined in (11).
@@
Next, we evaluate the impact of EVaR’s parameter $\alpha$ on the solution and policy of gambler’s ruin benchmark problem (Hau, Petrik, and Ghavamzadeh 2023; Bäuerle and Ott 2011). In the gambler’s ruin problem, a gambler starts with a given amount of capital and seeks to increase it up to some cap $K$. In each turn, the gambler decides how much of the available capital to stake. The stake doubles with a probability $q$ or is lost with a probability $1-q$. The gambler can also decide to quit and walk away with the current capital. The game also ends when the gambler either loses all capital or achieves (or surpasses) the cap $K$. The gambler collects the reward at the end, which equals the current capital. The exception is that the reward is -1 when the gambler loses all capital. The initial state is chosen uniformly among the non-terminal states.
The detailed description of the gambler’s ruin is as follows. The probability of winning a game is $q=0.68$, and the cap is $K=7$. The state space is $\bar{\mathcal{S}}=\{0,1,\cdots,7,e\}$. The action space is $\mathcal{A}=\{0,1,\cdots,6,X\}$. $X$ represents the action to quit the game. At state $s_{0}$, a gambler has capital 00, takes the only available action $a_{0}$, receives the reward $-1$, and transits to the sink state $e$. $\bar{p}(s_{0},a_{0},e)=1$ and $\bar{r}(s_{0},a_{0},e)=-1$. If the capital is $7$, a gambler reaches the target wealth level, takes the only available action $a_{0}$, receives the reward $7$, and transits to the sink state $e$. $\bar{p}(s_{7},a_{0},e)=1$ and $\bar{r}(s_{7},a_{0},e)=7$. For the sink state $e$, $\bar{p}(e,a_{0},e)=1$ and $\bar{r}(e,a_{0},e)=0$. For other states, in the $k$th game, the capital $c_{k}\in\{1,\cdots,6\}$, a gambler can take the action $a_{k}\in A(c_{k})=\{0,1,\cdots,c_{k}\}$ to continue $(k+1)$th game or take the action $X$ to quit the game. If $(k+1)$th game is won, $c_{(k+1)}=c_{k}+a_{k}$. If $(k+1)$th game is lost, $c_{k+1}=c_{k}-a_{k}$. Then we have $\bar{p}(s_{c_{k}},a_{k},s_{c_{k}+a_{k}})=q$, $\bar{r}(s_{c_{k}},a_{k},s_{c_{k}+a_{k}})=0$, $\bar{p}(s_{c_{k}},a_{k},s_{c_{k}-a_{k}})=1-q$, $\bar{r}(s_{c_{k}},a_{k},s_{c_{k}-a_{k}})=0$, $\bar{p}(s_{c_{k}},X,e)=1$, and $\bar{r}(s_{c_{k}},X,e)=c_{k}$. Given a risk level $\alpha$, we seek a policy that maximizes the EVaR of the random return defined in (14).
@@
Figure 3 shows optimal policies for 4 different EVaR risk levels $\alpha$ computed by Algorithm 1. The action indicates the amount of capital invested. The action 0 means quitting the game. Note that there is only one action when the capital is 00 and $7$ for all policies, so that action is neglected in Figure 3. Because the optimal policy is stationary, we can interpret and analyze it. The policies become notably less risk-averse as $\alpha$ increases. For example, when $\alpha=0.2$, the gambler is very risk-averse and always quits with the current capital. When $\alpha=0.4$, the gambler invests $1$ when capital is greater than $1$ and quits otherwise to avoid losing it all. When $\alpha=0.7$, the gambler always invests $1$ when capital is greater than 00. This causes the possibility of losing all capital. When $\alpha=0.9$, the gambler makes bigger bets, increasing the probabilities of reaching the cap and losing all capital.
To understand the impact of risk-aversion on the distribution of returns, we simulate the resulting policies over 7,000 episodes and show the distribution of capitals in Figure 4. When $\alpha=0.2$, the gambler takes a very conservative policy, and the return is a uniform distribution on the interval [1,7]. It is important to note that the horizon is a random variable and infinite. That is, if the gambler starts investing, the final capital will be boundary values. When $\alpha=0.4$, the gambler avoids the worst case of losing all capital by quitting with capital $1$, so the return is $1$ and $7$. When $\alpha=0.7$, if the capital is 1, the gambler still invests $1$. This increases the probability of reaching the cap. The return is $-1$ and $7$. When $\alpha=0.9$, the gambler makes bigger bets, and this increases the probability of losing all capital. The return is $-1$ and $7$. Overall, the figure shows that for lower values of $\alpha$, the gambler gives up some probability of reaching the cap in exchange for a lower probability of losing all capital.
6 Conclusion
We analyzed transient MDPs with two risk measures: ERM and EVaR. We established the necessary and sufficient conditions of the existence of stationary optimal polices. By removing the usual restriction that rewards are either all non-negative or all non-positive, we allow rewards that are negative as well as positive. We proved the convergence of value iteration and showed that the optimal value function is the unique solution to the exponential Bellman’s equation. We formulated linear program to compute the optimal exponential value functions. Our results illustrate the challenges of solving risk measure problems with TRC criterion and how our theoretical results effectively solve the problems. Future directions include extension to infinite-state TRC problems and risk-averse MDPs with average rewards. In this work, we assume that the states are fully observable, this work can also be extended with partial state observations.
Appendix A Background
@@
Tower Property for ERM For any two random variables $X_{1},X_{2}\in\mathbb{X}$, we have
$\operatorname{ERM}_{\beta}[X_{1}]=\operatorname{ERM}_{\beta}[\operatorname{ERM}_{\beta}[X_{1}|X_{2}]]$
@@
$\operatorname{ERM}_{\beta}[X_{1}|X_{2}]=-\beta^{-1}\log(\mathbb{E}[e^{-\beta X_{1}}|X_{2}])$
See, for example,  Hau, Petrik, and Ghavamzadeh (2023, theorem 3.1).
Appendix B Discounting versus TRC
@@
In the risk-neutral setting, it is well-known that the discounted objective can be interpreted as TRC (Altman 1998, Section 1.10). We summarize the construction in this section and show that it does not extend to risk-averse settings.
Given an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\mu)$, we construct an MDP ¯$\bar{\mathcal{M}}_{\gamma}=(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bar{\mu})$ such that $\bar{\mathcal{S}}=\mathcal{S}\cup\left\{e\right\}$, and $\bar{\mu}(s)=\mu(s),\forall s\in\mathcal{S}$ and $\mu(e)=0$. The transition function is defined as
$\bar{p}(s,a,s^{\prime})\;=\;\begin{cases}\gamma\cdot p(s,a,s^{\prime})\quad&\text{if}\quad s,s^{\prime}\in\mathcal{S},a\in\mathcal{A},\\ 1-\gamma\quad&\text{if}\quad s\in\mathcal{S},s^{\prime}=e,a\in\mathcal{A},\\ 1\quad&\text{if}\quad s=s^{\prime}=g,a\in\mathcal{A},\\ 0\quad&\text{otherwise}.\end{cases}$ (20)
When the rewards $r\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are independent of the next state, then $\bar{r}\colon\bar{\mathcal{S}}\times\bar{\mathcal{A}}\to\mathbb{R}$ are defined as $\bar{r}(s,a)=r(s,a)$ when $s\in\mathcal{S},a\in\mathcal{A}$ and $r(s,a)=0$ otherwise. The model can be readily extended to account for the target state dependence by constructing an ¯$\bar{\mathcal{M}}_{\gamma}$ with a random reward function. Any policy $\pi\in\Pi_{\mathrm{SD}}$ for $\mathcal{M}$ can be applied immediately to
@@
For each $\gamma\in[0,1)$ and MDP $\mathcal{M}$, and ¯$\bar{\mathcal{M}}_{\gamma}$ constructed above:
$\rho_{\gamma}(\mathcal{M},\pi,0)\;=\;g_{\infty}(\bar{\mathcal{M}}_{\gamma},\pi,0),\quad\pi\in\Pi_{\mathrm{SD}}.$
where the first parameters of $\rho$ and $g$ indicates the MDPs they apply to.
@@
The proposition follows from the construction above by algebraic manipulation (Altman 1998, section 1.10). ∎
@@
There exists an MDP $\mathcal{M}$, $\gamma>0$, $\beta>0$, and $\pi\in\Pi_{\mathrm{SD}}$ such that
$\rho_{\gamma}(\mathcal{M},\pi,0)\;\neq\;g_{\infty}(\bar{\mathcal{M}}_{\gamma},\pi,0),$
where the first parameters of $\rho$ and $g$ indicates the MDPs they apply to.
@@
Appendix C Theorems and Examples for Section 2
@@
Proof of Proposition C.1.
Consider the example of a transient MDP in Figure 1 with $r=-1$, a risk level $\alpha\in(0,1)$, and some $\epsilon\in(0,1)$ such that $\epsilon\geq\alpha$. This MDP admits only a single policy because it only has one action $a$ and we omit it in the notation below.
@@
$v_{t}(s)=\operatorname{CVaR}^{s}_{\alpha}\left[r(\tilde{s}_{0},a)+\operatorname{CVaR}_{\alpha}\left[r(\tilde{s}_{1},a)+\cdots\operatorname{CVaR}_{\alpha}\left[r(\tilde{s}_{t-1},a)\mid\tilde{s}_{t-2}\right]\dots\mid\tilde{s}_{1}\right]\right],$
The value function for the non-terminal state can be computed using a dynamic program as (Ahmadi et al. 2021a, theorem 1):
$v_{t}(s)$ $=\operatorname{CVaR}^{s}_{\alpha}[r(s,a)+v_{t-1}(\tilde{s}_{1})]$
$=-1+\min\left\{q_{1}\cdot v_{t-1}(s)\mid q\in\Delta^{2},\,q_{1}\leq\frac{\epsilon}{\alpha},q_{2}\leq\frac{1-\epsilon}{\alpha}\right\}$
@@ Appendix D Proofs for Section 3
$=-\frac{1}{\beta}\log\left(\cdot\sum_{k=0}^{t-1}(1-\epsilon)\epsilon^{k+1}\cdot\operatorname{exp}\left(-\beta\cdot r\right)^{k}\right).$
Here, $(1-\epsilon)\epsilon^{k}$ is the probability that the process terminates after exacly $k$ steps, and $\epsilon^{t}$ is the probability that the process does not terminate before reaching the horizon. Then, using the fact that a geometric series $\sum_{i=0}^{\infty}a\cdot q^{i}$ for $a\neq 0$ is bounded if and only if $|q|<1$ we get that
$\lim_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ >$\infty$> - ∞
[…]
$\epsilon\cdot\operatorname{exp}\left(-\beta\cdot r\right)$ $<1.$
@@ D.1 Optimality of Markov Policies
$\max_{\pi\in\Pi_{\mathrm{MR}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\max_{\pi\in\Pi_{\mathrm{HR}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$
See (Hau, Petrik, and Ghavamzadeh 2023, Corollary 4.2) for a proof. The result can also be derived from the optimality of Markov deterministic policies in MDPs with exponential utility functions (Chung and Sobel 1987; Patek 2001).

### v2

@@ 4.2 Algorithms
We now propose a simple algorithm for computing a $\delta$-optimal EVaR policy described in Algorithm 1. The algorithm reduces finding optimal EVaR-TRC policies to solving a sequence of ERM-TRC problems in (5). As Theorem 4.2 shows, there exists a $\delta$-optimal policy such that it is ERM-TRC optimal for some $\beta\in\mathcal{B}(\beta_{0},\delta)$. It is, therefore, sufficient to compute an ERM-TRC optimal policy for one of those $\beta$ values.
Data: MDP and desired precision $\delta>0$
@@
An additional advantage of Algorithm 1 is that the overhead of computing optimal solutions for multiple risk levels $\alpha$ can be small if one selects an appropriate set $\mathcal{B}$.
@@ 5 Numerical Evaluation
In this section, we illustrate our algorithms and formulations on tabular MDPs that include positive and negative rewards.
The ERM returns for the discounted and transient MDPs in Figure 1 with parameters $r=-0.2,\,\gamma=0.9,\,\epsilon=0.9$ are shown in Figure 2. The figure shows that, as expected, the returns are identical in the risk-neutral objective (when $\beta=0$). However, for $\beta>0$, the discounted and TRC returns differ significantly. The discounted return is unaffected by $\beta$ while the ERM-TRC return decreases with an increasing $\beta$. Please see Su et al. (2024a, appendix B) for more details.
@@
To evaluate the effect of risk-aversion on the structure of the optimal policy, we use the gambler’s ruin problem Hau et al. (2023b); Bäuerle and Ott (2011). In this problem, a gambler starts with a given amount of capital and seeks to increase it up to a cap $K$. In each turn, the gambler decides how much capital to bet. The bet doubles or is lost with a probability $q$ and $1-q$, respectively. The gambler can quit and keep the current wealth; the game also ends when the gambler goes broke or achieves the cap $K$. The reward equals the final capital, except it is -1 when the gambler is broke. The initial state is chosen uniformly. In the formulation, we use $q=0.68$, and a cap is $K=7$. The algorithm was implemented in Julia 1.10, and is available at https://github.com/suxh2019/ERMLP. Please see Su et al. (2024a, appendix F) for more details.
@@
Figure 3 shows optimal policies for four different EVaR risk levels $\alpha$ computed by Algorithm 1. The state represents how much capital the gambler holds. The optimal action indicates the amount of capital invested. The action 0 means quitting the game. Note that there is only one action when the capital is 00 and $7$ for all policies so that action is neglected in Figure 3. Because the optimal policy is stationary, we can interpret and analyze it. The policies become notably less risk-averse as $\alpha$ increases. For example, when $\alpha=0.2$, the gambler is very risk-averse and always quits with the current capital. When $\alpha=0.4$, the gambler invests $1$ when capital is greater than $1$ and quits otherwise to avoid losing it all. When $\alpha=0.9$, the gambler makes bigger bets, increasing the probability of reaching the cap and losing all capital.
@@
To understand the impact of risk-aversion on the distribution of returns, we simulate the resulting policies over 7,000 episodes and show the distribution of capitals in Figure 4. When $\alpha=0.2$, the return follows a uniform distribution on [1, 7]. When $\alpha=0.4$, the returns are $1$ and $7$. When $\alpha=0.7$ or $0.9$, the returns are $-1$ and $7$. Overall, the figure shows that for lower values of $\alpha$, the gambler gives up some probability of reaching the cap in exchange for a lower probability of losing all capital.
6 Conclusion and Future Work
We analyze transient MDPs with two risk measures: ERM and EVaR. We establish the existence of stationary deterministic optimal policies without any assumptions on the sign of the rewards, a significant departure from past work. Our results also provide algorithms based on value iteration, policy iteration, and linear programming for computing optimal policies.
Future directions include extensions to infinite-state TRC problems, risk-averse MDPs with average rewards, and partial-state observations.
Acknowledgments
We thank the anonymous reviewers for their detailed reviews and thoughtful comments, which significantly improved the paper’s clarity. This work was supported, in part, by NSF grants 2144601 and 2218063. Julien Grand-Clément was supported by Hi! Paris and Agence Nationale de la Recherche (Grant 11-LABX-0047).
A Background
@@
Tower Property for ERM For any two random variables $\tilde{x}_{1},\tilde{x}_{2}\in\mathbb{X}$, we have
$\operatorname{ERM}_{\beta}[\tilde{x}_{1}]=\operatorname{ERM}_{\beta}[\operatorname{ERM}_{\beta}[\tilde{x}_{1}\mid\tilde{x}_{2}]],$
@@
$\operatorname{ERM}_{\beta}[\tilde{x}_{1}|\tilde{x}_{2}]=-\beta^{-1}\log(\mathbb{E}[e^{-\beta\tilde{x}_{1}}\mid\tilde{x}_{2}])$
For a proof, see, for example, Hau et al. (2023b, theorem 3.1).
B Discounting versus TRC
@@
In the risk-neutral setting, it is well-known that the discounted objective can be interpreted as TRC (Altman, 1998, Section 1.10). We summarize the construction in this section and show that it does not extend to risk-averse settings.
Given an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\mu)$, we construct a transient MDP ¯$\bar{\mathcal{M}}_{\gamma}=(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bar{\mu})$ such that $\bar{\mathcal{S}}=\mathcal{S}\cup\left\{e\right\}$, and $\bar{\mu}(s)=\mu(s),\forall s\in\mathcal{S}$ and $\mu(e)=0$. The transition function $\bar{p}$ is defined as
$\bar{p}(s,a,s^{\prime})\;=\;\begin{cases}\gamma\cdot p(s,a,s^{\prime})\quad&\text{if}\quad s,s^{\prime}\in\mathcal{S},a\in\mathcal{A},\\ 1-\gamma\quad&\text{if}\quad s\in\mathcal{S},s^{\prime}=e,a\in\mathcal{A},\\ 1\quad&\text{if}\quad s=s^{\prime}=e,a\in\mathcal{A},\\ 0\quad&\text{otherwise}.\end{cases}$ (21)
When the rewards $r\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are independent of the next state, then $\bar{r}\colon\bar{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ are defined as
$\bar{r}(s,a)=\begin{cases}r(s,a)&\text{if }s\in\mathcal{S},\\ 0&\text{otherwise}.\end{cases}$ (22)
The model can be readily extended to account for the target state dependence by constructing an ¯$\bar{\mathcal{M}}_{\gamma}$ with the reward function $\bar{r}\colon\bar{\mathcal{S}}\times\mathcal{A}\times\bar{\mathcal{S}}\to\mathbb{R}$ as
$\bar{r}(s,a,s^{\prime})=\begin{cases}\gamma^{-1}\cdot r(s,a,s^{\prime})&\text{if }s,s^{\prime}\in\mathcal{S},\\ 0&\text{otherwise}.\end{cases}$
Recall that when $\beta=0$, we have $\lim_{\beta\rightarrow 0^{+}}\operatorname{ERM}[\tilde{x}]=\mathbb{E}[\tilde{x}]$.
Proposition B.1 shows that the expected total returns are identical in a discounted MDP $\mathcal{M}$ and in a transient MDP ¯$\bar{\mathcal{M}}_{\gamma}$.
@@
For each $\gamma\in[0,1)$ and MDPs ,¯$\mathcal{M},\bar{\mathcal{M}}_{\gamma}$ constructed above, we have that
$\rho_{\gamma}(\pi,0)\;=\;g_{\infty}(\pi,0),\quad\pi\in\Pi_{\mathrm{SD}}.$
where $\rho_{\gamma}(\pi,0)$ is defined in (20) and $g_{\infty}(\pi,0)$ is defined in (6).
@@
The proposition follows from the construction above by algebraic manipulation (Altman, 1998, section 1.10). ∎
@@
There exists an MDP $\mathcal{M}$, $\gamma>0$, $\beta>0$, ¯$\bar{\mathcal{M}}_{\gamma}$ constructed above, and $\pi\in\Pi_{\mathrm{SD}}$ such that
$\rho_{\gamma}(\pi,\beta)\;\neq\;g_{\infty}(\pi,\beta),$
where $\rho_{\gamma}(\pi,0)$ is defined in (20) and $g_{\infty}(\pi,0)$ is defined in (6).
@@
C Additional Results of Section 2
@@
Proof.
Consider the example of a transient MDP in Figure 1 with $r=-1$, a risk level $\alpha\in(0,1)$, and some $\epsilon\in(0,1)$ such that $\epsilon\geq\alpha$. This MDP admits only a single policy because it only has one action $a$, which we omit in the notation below.
@@
$v_{t}(s)\;=\;\operatorname{CVaR}^{s}_{\alpha}\left[r(\tilde{s}_{0},a)+\operatorname{CVaR}_{\alpha}^{\tilde{s}_{1}}\left[r(\tilde{s}_{1},a)+\cdots\operatorname{CVaR}^{\tilde{s}_{t-1}}_{\alpha}\left[r(\tilde{s}_{t-1},a)\right]\right]\right],$
The value function for the non-terminal state can be computed using a dynamic program as (Ahmadi et al., 2021a, theorem 1):
$v_{t}(s)$ $=\operatorname{CVaR}^{s}_{\alpha}[r(s,a)+v_{t-1}(\tilde{s}_{1})]=-1+\min\left\{q_{1}\cdot v_{t-1}(s)\mid q\in\Delta^{2},\,q_{1}\leq\frac{\epsilon}{\alpha},q_{2}\leq\frac{1-\epsilon}{\alpha}\right\}$
@@ D Proofs of Section 3
$=-\frac{1}{\beta}\log\left(\sum_{k=0}^{t-1}(1-\epsilon)\epsilon^{k}\cdot\operatorname{exp}\left(-\beta\cdot r\right)^{k}\right).$
Here, $(1-\epsilon)\epsilon^{k}$ is the probability that the process terminates after exactly $k$ steps, and $\epsilon^{t}$ is the probability that the process does not terminate before reaching the horizon. Then, using the fact that a geometric series $\sum_{i=0}^{\infty}a\cdot q^{i}$ for $a\neq 0$ is bounded if and only if $|q|<1$ we get that
$\lim_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]>-\infty\qquad\Leftrightarrow\qquad\epsilon\cdot\operatorname{exp}\left(-\beta\cdot r\right)<1.$
@@ D.1 Optimality of Markov Policies
$\begin{split}\max_{\pi\in\Pi_{\mathrm{MD}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\max_{\pi\in\Pi_{\mathrm{HR}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].\end{split}$
See (Hau et al., 2023b, Corollary 4.2) for a proof. The result can also be derived from the optimality of Markov deterministic policies in MDPs with exponential utility functions Chung and Sobel (1987); Patek (2001).